A box with mass 1 kg slides down an inclined plane for a distance of 1m. The inclined plane makes an angle of 30° with the grFrom the choices given below. choose the correct magnitude and direction of the center of mass of the system of particles shoMagnitude: approx. 1.85 m Direction: 45° from X axis direction Magnitude: approx. 0.85 m Direction: 45° from X axis directionA1500-kg elevator car moves 40.0 m above the ground at constant speed. (a) What is the work done on the lift by the gravitatiI need all 3 questions please
Answer #1

1.

work done is given by,

W = F*d*cos(A)

here, F = force = friction force

F = \mu *N

d = distance travelled = 1 m

A = angle between F & d = 180 deg

where,\mu = 0.5

N = normal force = m*(g*cos\theta)

given \theta = inclined angle = 30 deg

m = mass = 1 kg

So, N = 1*9.8*cos(30 deg) = 8.5

then, W = 0.5*8.5*1*cos(180 deg)

W = -4.25 J

therefore correct option is B. Approximately -4.2 J

2.

Center of mass is given by:

In x-direction

Xcm = (m1x1 + m2x2 + m3x3)/(m1 + m2 + m3)

In y-direction

Ycm = (m1y1 + m2y2 + m3y3)/(m1 + m2 + m3)

m1 = 2 kg, (x1, y1) = (0,0)

m2 = 2 kg, (x2, y2) = (1, 0)

m3 = 2 kg, (x3, y3) = (0, 1)

m4 = 4 kg, (x4, y4) = (1, 1)

Now

Xcm = (2*0 + 2*1 + 2*0 + 4*1)/(2+2+2+4) = 0.6 m

Ycm = (2*0 + 2*0 + 2*1 + 4*1)/(2+2+2+4) = 0.6 m

Center of mass will be at (Xcm, Ycm) = (0.6, 0.6) m

So, magnitude of distance of center of mass from origin = d = sqrt(0.6^2 + 0.6^2)

d = 0.8485 m = approx. 0.85 m

direction of center of mass = \theta = arctan(Ycm/Xcm) = arctan(1/1)

\theta = 45 deg from X axis direction

therefore correct option is B.

Magnitude : approx. 0.85 m

Direction : 45 deg from X axis direction

Please upvote.